OpenStudy (anonymous):

David uses 10 lemons to make 1 cup of lemonade. How many lemons would he use to make 1 pint? a. 40 lemons c. 20 lemons b. 5 lemons d. 10 lemons

OpenStudy (texaschic101):

1 pint = 2 cups 10/1 = x/2 (10 lemons to 1 cup = x lemons to 2 cups) cross multiply (1)(x) = (10)(2) x = 20 lemons

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

The Natchez Trace Bridge in Franklin, Tennessee, is 1,500 feet long. Suppose you build a model of the bridge using a scale of 1 inch:25 feet. How long is the model?

OpenStudy (anonymous):

what about this one

OpenStudy (texaschic101):

1500/25 = 60 inches.

OpenStudy (anonymous):

nope not there in my answers box

OpenStudy (anonymous):

but thank yoou any way

OpenStudy (texaschic101):

let me try again...hold on

OpenStudy (texaschic101):

1/25 = x/1500 (1 inch to 25 ft = x inches to 1500 ft) cross multiply (25)(x) = (1)(1500) 25x = 1500 x = 1500/25 x = 60 inches I am getting the same answer....what are your answer choices....are they in feet or inches ?

OpenStudy (anonymous):

here the answersa. 1:57 c. 1 in.:65 ft b. 1 in.:61 ft d. 1:61

OpenStudy (texaschic101):

hold on...lets get some help @whpalmer4 ....this is not history...it is just in the wrong subject...its math

OpenStudy (anonymous):

yep

OpenStudy (texaschic101):

@amistre64 .....can you take a peek at this please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what is it

OpenStudy (texaschic101):

@skullpatrol .....can you help

OpenStudy (texaschic101):

you wouldn't have to wait for help if this was in the math section....after all, this is math, not history

OpenStudy (texaschic101):

@primeralph ...help

OpenStudy (primeralph):

How many cups = 1 pint?

OpenStudy (texaschic101):

no..not the lemon one...the next question

OpenStudy (primeralph):